OpenStudy (anonymous):

Saifoo.khan please help with the sqrt(240)

OpenStudy (anonymous):

\[240=16\times 15\] is a hint

OpenStudy (anonymous):

\[16*\sqrt{15}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

\[\sqrt{240}=4\sqrt{15}\]

OpenStudy (anonymous):

safiah

OpenStudy (anonymous):

when you break it down and there is two groups of 2 do you just do 2(2)?

OpenStudy (asnaseer):

@satellite73 gave you a very big hint at ow to solve this. you need to make use of the rule \(\sqrt{a\times b}=\sqrt{a}\times\sqrt{b}\). so, in your case, we have \(\sqrt{240}=\sqrt{16\times15}=\sqrt{16}\times\sqrt{15}=4\times\sqrt{15}=4\sqrt{15}\)